An investor has $100,000 to invest. The investor faces two choices: buying physical gas or gas call options. The call option with the $3.25 strike price is currently selling at $0.25, with the current market price of gas being 3.45.

  1. What’s reason behind the investor’s decision to buy (either gas or the call)?
  2. How high does the gas price have to rise for the option strategy to be more profitable?
  3. What’s your recommendation to the investor?
Answer #1
a Reason behind decision to buy:
Expectation of increase in price of gas
If price of gas goes up, there will be profit by selling at higher price
b Price of call option $0.25
Number of call option that can be purchased          400,000 ($100000/$0.25)
Market price of gas per unit $3.45
Number of units of gas can be purchased            28,986 (100000/3.45)
Payoff on calloption =400000*(P-3.25)
P=Price of gas at settlement
Profit on purchase of gas=28986*(P-3.45)
P A=400000*(P-3.25) B=A-100000 C=28986*(P-3.45)
Price of gas at settlement Payoff on call option Profit on Call Option Profit on purchase of gas
$0.05 $0 ($100,000) ($98,552.40)
$1.00 $0 ($100,000) ($71,015.70)
$1.50 $0 ($100,000) ($56,522.70)
$2.00 $0 ($100,000) ($42,029.70)
$2.50 $0 ($100,000) ($27,536.70)
$3.00 $0 ($100,000) ($13,043.70)
$3.25 $0 ($100,000) ($5,797.20)
$3.30 $20,000 ($80,000) ($4,347.90)
$3.35 $40,000 ($60,000) ($2,898.60)
$3.40 $60,000 ($40,000) ($1,449.30)
$3.45 $80,000 ($20,000) ($0.00)
$3.50 $100,000 ($0) $1,449.30
$3.51 $104,000 $4,000 $1,739.16
$3.55 $120,000 $20,000 $2,898.60
$3.60 $140,000 $40,000 $4,347.90
$3.65 $160,000 $60,000 $5,797.20
Option Strategy will be more profitable if the price rises to $3.51
c RECOMMENDATION TO INVESTOR:
Option Strategy is High Risk high gain strategy
If prices falls below $3.25, $100000 will be lost
Gain will be significantly high if price goes up
Buy strategy will give limited gain and limited loss
Option Strategy should be chosen if the investor is reasonably sure that price willnot fall
Investor with limited risk preference should choose BUY Strategy